One number is 8 more than another number. The sum of the number is 12 find the 2 numbers
Answer by ikleyn(52921)

.
x + (x+8) = 12


2x + 8 = 12


2x = 12 - 8 = 4  ====>  x = 4/2 = 2.


Answer.  The numbers are  2  and  2+8 = 10.

Solved.